Office Depot sells boxes of mechanical pencils ($10) and paper clips ($4). Tsukiko ordered a total of 24 boxes for $210. How many boxes of each did Tsukiko order?
Answer #1

Let's Assume x = Pencil boxes & y = paper clips.

x($10) + y($4) = $210

10x + 4y = 210 ------------------------ (1)

x + y = 24 boxes

x = 24-y ---------------------- (2)

Insert equation (2) in equation (1)

10 (24-y) + 4y = 210

240-10y+4y = 210

240-6y=210

6y=240-210

6y=30

y=30/6

y=5

x = 24-5 = 19

Therefore, Tsukiko ordered Pencil boxes(x) are 19 boxes and Paper clip boxes(y) are 5